OpenStudy (anonymous):

What is the simplified form of x/2/4/2x

OpenStudy (anonymous):



OpenStudy (phi):

when you divide by a fraction, you can "flip" the fraction and then multiply what is 4/(2x) flipped ?

OpenStudy (anonymous):

2x/4

OpenStudy (phi):

yes. so instead of dividing by 4/(2x), multiply by 2x/4 \[ \frac{x}{2} \cdot \frac{2x}{4} \] I would first simplify, by noticing that 2/2 is 1 (so "cancel" a 2 from the top and bottom) then multiply top times top, and bottom times bottom

OpenStudy (anonymous):

The answer is X^2 / 4

OpenStudy (mathmate):

Also, when you have a "superimposed fraction" like this, I would calculate \(\Large\frac{outer}{inner}\). For example: \(\huge\frac{\frac{6}{5}}{\frac{24}{15}}=\frac{6*15}{5*24}\) DO NOT multiply until you have cancelled all common factors.

OpenStudy (mathmate):

Yes, answer is correct.